# NeuroIDBench: An Open-Source Benchmark Framework for the Standardization of Methodology in Brainwave-based Authentication Research

### v3

@@ 1 Introduction
In today’s rapidly changing security landscape, the authentication process is a crucial element of access control mohamed2022systematic . The increasing impracticality and safety concerns associated with conventional authentication methods, such as passwords alroomi2023measuring ; taneski2014password , are driving the development of novel biometric solutions. In this domain, user recognition based on brain activity has gained increasing attention gui2019survey , especially with advancements in consumer-grade EEG (Electroencephalogram) technology emotivEPOCChannel ; Neurable ; Muse ; Varjo . A notable application scenario is Extended Reality (XR), where seamless and secure biometric authentication systems are urgently needed stephenson2022sok . Besides, brain biometrics bring broader benefits, such as enabling hands-free authentication, resistance to observation, inherent liveness detection, broad universality, and potential for revocability gui2019survey ; zhang2021review ; lin2018brain .
@@
Unfortunately, the research in brainwave-based authentication is beset by problems of generalizability, comparability, and reproducibility. To begin, dataset sharing is sharply curtailed by ethical considerations and privacy laws. Therefore, many studies are based on self-collected undisclosed datasets, which often leads to results that are potentially overfitted or optimized to the specific characteristics of those datasets, with uncertain applicability to other, independent datasets. Also, the size of the datasets used in studies is severely limited by the complexity and resource intensiveness of broad data collection. In fact, most brainwave-based authentication studies have much fewer than 50 subjects gui2019survey ; bidgoly2020survey , so that the evaluations are normally conducted on very small datasets. Consequently, the models and hyperparameters are prone to overfitting, which means the reported results will likely fail to generalize. To make matters worse, too many authors either neglect making available their source code or provide code which is inoperable, with the consequence that researchers are prevented from reproducing and building upon previous work. Clearly, under circumstances such as these, our subcommunity of authentication researchers will not make any significant progress. That is why the time has come for us to consolidate our practice and advance the research toward consistent, reproducible, and generalizable findings. To address this gap, we make the following contributions:
@@
  • (1) We present NeuroIDBench, the first comprehensive open-source [^1] https://github.com/Avichaurasia/NeuroIDBench.git benchmark tool designed to assist researchers in evaluating their brainwave-based authentication approaches. NeuroIDBench is engineered to allow for the easy and flexible integration of new datasets and methods, enabling fair comparability. NeuroIDBench incorporates nine public datasets, which together contain a substantial number of subjects (n = 285, Table 2) and multi-session recordings. To determine the impact of preprocessing on performance, NeuroIDBench allows to apply different sample rejection thresholds as well as sample length parameters (Sec 4.1) and feature extraction methods (Sec 4.2). Also, it implements a baseline with the most popular authentication algorithms in the state of the art, including shallow classifiers and deep learning-based approaches. Lastly, it incorporates two attack scenarios under which to test the algorithms mansfield2002best : i) known attacker, where the attacker is known to the authentication model (i.e., enrolled in the system), and, ii) unknown attacker, a more realistic situation in which the attacker biometric data has not been seen by the system previous to the attack.
@@
Among the key findings in the analysis we observe that Random Forest consistently outperforms other classifiers and its performance is comparable to that of Twin Neural Networks. As the latter brings not only better performance but scalability advantages, we recommend more research on deep learning-based approaches to move forward, which is also the path followed in other biometric communities wang2021deep ; rim2021fingerprint ; minaee2023biometrics . Furthermore, our results confirm that integrating Power Spectral Density (PSD) with an Autoregressive (AR) model of order 1 stands out as the most effective feature extraction approach. When investigating the impact of the adversary model, our results indicate that the average Equal Error Rate (EER) for known attackers is 2.87%, while for unknown attackers, it is 4.6%. This trend in performance degradation underscores the need to evaluate brainwave authentication solutions under unknown attacker scenarios to provide a realistic performance indicator, a practice that is uncommon in current work. Variations in the performance across datasets confirm that brainwave-based authentication algorithms validated on single datasets are not representative, highlighting the need for more robust validation in the community. Finally, the observed disparities in the results for single-session versus multi-session authentication reveal a substantial gap that requires future research in this area.
@@ 2.1 Using brainwaves for Biometric Authentication
Brainwaves, or EEG signals, represent the electrical activity produced by neuron interactions in the brain niedermeyer2005electroencephalography ; hu2019eeg . These signals are commonly recorded from the scalp using specialized sensors. Originally developed for medical purposes haas2003hans , such as diagnosing neurological disorders adeli2010automated , EEG technology has expanded its applications significantly. Its usage now extends to Brain-Computer Interfaces (BCI), facilitating direct interaction between the human brain and external devices kawala2021summary , and enjoys acceptance in the consumer market, especially for gaming and health-related applications.
Identity verification, commonly known as authentication, involves a one-to-one comparison to confirm if an individual presenting a previously registered biometric trait is indeed the same person. In essence, it addresses the question: ”Are you who you claim to be?” This process is crucial to differentiate legitimate users from imposters. Biometric authentication comprises two primary phases: enrollment and verification ISO2382-37 . During enrollment, users’ unique biometric characteristics are captured and stored as templates or used to train and store a classification model for that user. Each template/user model is linked to a unique user identifier, such as a username. In the verification phase, the system compares the current biometric sample with the stored template or runs it through the classification model, corresponding to the claimed identity.
Figure 1: Overview of elements in a brainwave-based biometric authentication process  arias2023performance .
@@
    Raw EEG Data Acquisition: This initial step captures the brain’s electrical activity. An effective way to elicit signals useful for authentication is measuring brain reactions to controlled stimuli, such as audio or images. These time-locked reactions,called Event-Related Potentials (ERPs), have shown unique features to individuate people arias2021inexpensive and are distinguished by their high Signal-to-Noise Ratio (SNR) armstrong2015brainprint ; zhang2021review . Moreover, the ability to modify the stimuli associated with ERPs enhances the system’s revocability. This feature is invaluable in mitigating risks associated with potential data compromise lin2018brain . NeuroIDBench focuses in ERP-based datasets, given their superior benefits with regard to alternative brainwave elicitation paradigms.
@@
    Signal Pre-processing: The acquired raw EEG data is then refined through pre-processing. This may include filtering out noise, rejecting artifacts, and downsampling the signal, aiming to isolate the purest form of the EEG data for authentication. purposes gui2019survey . NeuroIDBench embeds a pre-processing pipeline based on best-practices for handling ERPs as used in the community gui2019survey .
@@ 2.2 Adversary Model
In addressing the security of brainwave-based authentication systems, we consider the zero-effort attacker as the primary threat model. This approach aligns with the standard threat model used in most brainwave authentication studies. A zero-effort attack is characterized by an adversary who attempts to gain unauthorized access using their own biometric data, without employing any advanced techniques to mimic or falsify another individual’s biometric trait mansfield2002best . In this context, the key challenge in protecting against zero-effort attacks lies in the biometric system’s ability to accurately distinguish between the biometric data of an actual user and that of the intruder. The system must be adept at minimizing false matches, where the adversary’s biometric data is erroneously accepted as that of a legitimate user.
In our consideration of zero-effort attacks, we distinguish between two types of attackers: known and unknown. A known attacker refers to one whose biometric data has previously been exposed to the system, typically during the training phase. This exposure can influence the system’s learning and adaptation, potentially affecting its ability to accurately identify or reject these attackers later. On the other hand, an unknown attacker is one whose biometric data has never been introduced to the system. This lack of prior exposure means that the system has no learning or adaptation based on this attacker’s data, making it a more realistic and challenging scenario. We expect that unknown attackers represent a higher risk to the system, as they test the system’s ability to authenticate users based purely on learned patterns without prior knowledge of specific individuals. Our focus on defending against unknown attackers is driven by the need to develop robust authentication systems that can reliably identify and prevent unauthorized access by individuals completely unknown to the system.
Figure 2: Schematic depiction outlining the architectural structure of the NeuroIDBench codebase, providing a visual overview of its underlying framework jayaram2018moabb
@@ 2.3 Related Work
Arias et al.(2023) arias2023performance 52,40 $\times$
Fallahi et al.(2023) fallahi2023brainnet 41,40
Hernandez et al.(2022) hernandez2022eeg 39 $\times$ $\times$
Bidgoly et al.(2022) bidgoly2022towards 109 $\times$
Maiorana(2021) maiorana2021learning 45 $\times$ $\times$
Sooriyaarachchi et al.(2020) sooriyaarachchi2020musicid 20 $\times$ $\times$
Gupta et al.(2020) gupta2020blink 20
Nakanishi et al.(2019) nakanishi2019biometric 10 $\times$ $\times$
Lin et al.(2018) lin2018brain 179 $\times$ $\times$
Schons et al.(2018) schons2018convolutional 109 $\times$
Das et al.(2016) das2016eeg 50 $\times$ $\times$
@@
Presently, two primary methods for brainwave authentication are studied in the literature. The first method utilizes shallow classifiers built on time series-based feature extraction techniques, such as Power Spectral Density (PSD) or Autoregressive (AR) models arias2023performance ; arias2021inexpensive ; nakanishi2019biometric ; das2016eeg . The second approach involves training specialized deep learning models designed explicitly for brainwave data, aiming to extract features directly associated with user identity fallahi2023brainnet ; maiorana2021learning ; schons2018convolutional ; bidgoly2022towards . However, traditional feature extraction methods are not sufficiently effective, as they were not originally designed for authentication purposes. Meanwhile, deep learning approaches require a substantial amount of data to develop an adequate feature extractor, which is currently scarce in the field of brainwave authentication. Therefore, it is crucial to determine the most appropriate hyperparameters based on other datasets in the field, which can aid in mitigating these challenges.
@@
Table 1 presents an analysis of 11 papers in the realm of brainwave authentication. Among these, only two papers have made their source code available fallahi2023brainnet ; gupta2020blink , and merely five have utilized public datasets fallahi2023brainnet ; gupta2020blink ; arias2023performance ; bidgoly2022towards ; schons2018convolutional . Notably, four of these papers employed two common public datasets erpcore ; schalk2004bci2000 , reflecting the limited data availability, which consequently diminishes the generalizability of the results. On average, these studies included approximately 70 subjects for their evaluation purposes. The subcommunity in brainwave-based authentication clearly require more transparency in the research practice as well as more accessibility in the datasets and source code. To be sure, the brainwave-authentication subcommunity are, in this regard, not alone. For example, Olszewski et al. olszewski2023get find that the ”large and long-lived” subcommunity of machine learning security fail to provide data (30%) or source code (40%) even in the papers published at the top-tier security conferences. The authors found out that a full 80% of artifacts are incapable of reproducing the reported results. It is therefore unsurprising that similar problems should plague the new and emerging subcommunity in brainwave-based authentication. Nevertheless, these issues are intensified in brainwave authentication due to the relatively small size of each dataset, heightening the risk of overfitting.
@@
Meanwhile, researchers in the field of Brain-Computer Interface (BCI) encounter similar challenges. Hence, Jayaram and Barachant moabb proposed a benchmark framework [^2] https://github.com/NeuroTechX/moabb for investigating common BCI algorithms across 22 public EEG datasets with over 250 participants. This framework aims to address issues of reproducibility, lack of open-source resources, and result compatibility. However, their study did not cover authentication algorithms. Therefore, it is vital for brainwave authentication research to adopt an open-source approach and to extensively employ existing public datasets. It is crucial to take such steps to enhance the robustness of research in this field and to promote a deeper, more comprehensive exploration of brainwave authentication techniques.
@@
N400 Pijnacker et al. pijnacker2017semantic 2017 45 actiCap 32 500 Hz 1
N400 Draschkow et al. draschkow2018no 2018 40 BrainAmp, actiChamp 64 1000 Hz 1
N400 Marzecová et al. marzecova2018attentional 2018 18 BrainAmp 59 500 Hz 1
N400 Mantegna et al. mantegna2019distinguishing 2019 31 BrainAmp, EasyCap 65 500 Hz 1
N400 ERPCORE: N400 erpcore 2021 40 Biosemi 30 1024 Hz 1
N400 Hodapp and Rabovky hodapp2021n400 2021 33 BrainAmp 64 1000 Hz 1
N400 Rabs et al. rabs2022situational 2022 38 BrainVision 26 500 Hz 1
N400 Schoknecht et al. schoknecht2022interaction 2022 38 ActiCap, ActiChamp 58 500 Hz 1
N400 Toffolo et al. toffolo2022evoking 2022 24 Biosemi 128 512 Hz 1
N400 Lindborg et al. lindborg2023semantic 2022 40 BrainVision 64 2046 Hz 1
N400 COGBCI: Flanker cogbci 2023 29 ActiCap, actiChamp 64 512 Hz 3
N400 Stone et al. stone2023understanding 2023 64 TMSi Refa 32 512 Hz 1
P300 BrainInvaders12 van2019building 2012 25 NeXus-32 16 128 Hz 1
P300 BrainInvaders13a vaineau2019brain 2013 24 g.GAMMAcap 16 512 Hz 1
P300 BrainInvaders14a bi2014a 2014 64 g.Sahara 16 512 Hz 1
P300 BrainInvaders14b bi2014b 2014 37 g.GAMMAcap 32 512 Hz 1
P300 Gao et al. gao2014novel 2014 30 Neuroscan 12 500 Hz 1
P300 BrainInvaders15a brainInvaders15a 2015 50 g.GAMMAcap 32 512 Hz 1
P300 BrainInvaders15b bi2015b 2015 44 g.GAMMAcap 32 512 Hz 1
P300 Mouček et al. mouvcek2017event 2017 250 BrainVision 3 n.a. 1
P300 Hubner et al. hubner2017learning 2017 13 BrainAmp DC 31 1000 Hz 1
P300 Sosulski and Tangermann sosulski2019spatial 2019 13 BrainAmp, EasyCap 31 1000 Hz 1
P300 Lee et al. lee2019eeg 2019 54 BrainAmp 62 1000 Hz 2
P300 Simões et al. simoes2020bciaut 2020 15 g.tec 8 250 Hz 7
P300 Goncharenko et al. goncharenko2020raccoons 2020 60 NVX-52 8 500 Hz 1
P300 Chatroudi et al. houshmand2021effect 2021 24 g.tec 64 1200 Hz 1
P300 Cattan et al. cunningham2021underestimation 2021 21 g.USBamp, g.tec 16 512 Hz 1
P300 ERPCORE: P300 erpcore 2021 40 Biosemi 30 1024 Hz 1
P300 Won et al. won2022eeg 2022 55 Biosemi 32 512 Hz 1

### v5

@@ 1 Introduction
In today’s rapidly changing security landscape, the authentication process is a crucial element of access control [1]. The increasing impracticality and safety concerns associated with conventional authentication methods, such as passwords [2, 3], are driving the development of novel biometric solutions. In this domain, user recognition based on brain activity has gained increasing attention [4], especially with advancements in consumer-grade EEG (Electroencephalogram) technology [5, 6, 7, 8]. A notable application scenario is Extended Reality (XR), where seamless and secure biometric authentication systems are urgently needed [9]. Besides, brain biometrics bring broader benefits, such as enabling hands-free authentication, resistance to observation, inherent liveness detection, broad universality, and potential for revocability [4, 10, 11].
@@
Unfortunately, the research in brainwave-based authentication is beset by problems of generalizability, comparability, and reproducibility. To begin, dataset sharing is sharply curtailed by ethical considerations and privacy laws. Therefore, many studies are based on self-collected undisclosed datasets, which often leads to results that are potentially overfitted or optimized to the specific characteristics of those datasets, with uncertain applicability to other, independent datasets. Also, the size of the datasets used in studies is severely limited by the complexity and resource intensiveness of broad data collection. In fact, most brainwave-based authentication studies have much fewer than 50 subjects [4, 12], so that the evaluations are normally conducted on very small datasets. Consequently, the models and hyperparameters are prone to overfitting, which means the reported results will likely fail to generalize. To make matters worse, too many authors either neglect making available their source code or provide code which is inoperable, with the consequence that researchers are prevented from reproducing and building upon previous work. Clearly, under circumstances such as these, our subcommunity of authentication researchers will not make any significant progress. That is why the time has come for us to consolidate our practice and advance the research toward consistent, reproducible, and generalizable findings. To address this gap, we make the following contributions:
@@
  • (1) We present NeuroIDBench, the first comprehensive open-source [^1] https://github.com/Avichaurasia/NeuroIDBench.git benchmark tool designed to assist researchers in evaluating their brainwave-based authentication approaches. NeuroIDBench is engineered to allow for the easy and flexible integration of new datasets and methods, enabling fair comparability. NeuroIDBench incorporates nine public datasets, which together contain a substantial number of subjects (n = 285, Table 2) and multi-session recordings. To determine the impact of preprocessing on performance, NeuroIDBench allows to apply different sample rejection thresholds as well as sample length parameters (Sec 4.1) and feature extraction methods (Sec 4.2). Also, it implements a baseline with the most popular authentication algorithms in the state of the art, including shallow classifiers and deep learning-based approaches. Lastly, it incorporates two attack scenarios under which to test the algorithms [13]: i) known attacker, where the attacker is known to the authentication model (i.e., enrolled in the system), and, ii) unknown attacker, a more realistic situation in which the attacker biometric data has not been seen by the system previous to the attack.
@@
Among the key findings in the analysis we observe that Random Forest consistently outperforms other classifiers and its performance is comparable to that of Twin Neural Networks. As the latter brings not only better performance but scalability advantages, we recommend more research on deep learning-based approaches to move forward, which is also the path followed in other biometric communities [14, 15, 16]. Furthermore, our results confirm that integrating Power Spectral Density (PSD) with an Autoregressive (AR) model of order 1 stands out as the most effective feature extraction approach. When investigating the impact of the adversary model, our results indicate that the average Equal Error Rate (EER) for known attackers is 2.87%, while for unknown attackers, it is 4.6%. This trend in performance degradation underscores the need to evaluate brainwave authentication solutions under unknown attacker scenarios to provide a realistic performance indicator, a practice that is uncommon in current work. Variations in the performance across datasets confirm that brainwave-based authentication algorithms validated on single datasets are not representative, highlighting the need for more robust validation in the community. Finally, the observed disparities in the results for single-session versus multi-session authentication reveal a substantial gap that requires future research in this area.
@@ 2.1 Using brainwaves for Biometric Authentication
Brainwaves, or EEG signals, represent the electrical activity produced by neuron interactions in the brain [17, 18]. These signals are commonly recorded from the scalp using specialized sensors. Originally developed for medical purposes [19], such as diagnosing neurological disorders [20], EEG technology has expanded its applications significantly. Its usage now extends to Brain-Computer Interfaces (BCI), facilitating direct interaction between the human brain and external devices [21], and enjoys acceptance in the consumer market, especially for gaming and health-related applications.
Identity verification, commonly known as authentication, involves a one-to-one comparison to confirm if an individual presenting a previously registered biometric trait is indeed the same person. In essence, it addresses the question: ”Are you who you claim to be?” This process is crucial to differentiate legitimate users from imposters. Biometric authentication comprises two primary phases: enrollment and verification [22]. During enrollment, users’ unique biometric characteristics are captured and stored as templates or used to train and store a classification model for that user. Each template/user model is linked to a unique user identifier, such as a username. In the verification phase, the system compares the current biometric sample with the stored template or runs it through the classification model, corresponding to the claimed identity.
Figure 1: Overview of elements in a brainwave-based biometric authentication process  [23].
@@
    Raw EEG Data Acquisition: This initial step captures the brain’s electrical activity. An effective way to elicit signals useful for authentication is measuring brain reactions to controlled stimuli, such as audio or images. These time-locked reactions, called Event-Related Potentials (ERPs), have shown unique features to individuate people [24] and are distinguished by their high Signal-to-Noise Ratio (SNR) [25, 10]. Moreover, the ability to modify the stimuli associated with ERPs enhances the system’s revocability. This feature is invaluable in mitigating risks associated with potential data compromise [11]. NeuroIDBench focuses in ERP-based datasets, given their superior benefits with regard to alternative brainwave elicitation paradigms.
@@
    Signal Pre-processing: The acquired raw EEG data is then refined through pre-processing. This may include filtering out noise, rejecting artifacts, and downsampling the signal, aiming to isolate the purest form of the EEG data for authentication purposes [4]. NeuroIDBench embeds a pre-processing pipeline based on best-practices for handling ERPs as used in the community [4].
@@ 2.2 Adversary Model
In addressing the security of brainwave-based authentication systems, we consider the zero-effort attacker as the primary threat model. This approach aligns with the standard threat model used in most brainwave authentication studies. A zero-effort attack is characterized by an adversary who attempts to gain unauthorized access using their own biometric data, without employing any advanced techniques to mimic or falsify another individual’s biometric trait [13]. In this context, the key challenge in protecting against zero-effort attacks lies in the biometric system’s ability to accurately distinguish between the biometric data of an actual user and that of the intruder. The system must be adept at minimizing false matches, where the adversary’s biometric data is erroneously accepted as that of a legitimate user.
In our consideration of zero-effort attacks, we distinguish between two types of attackers: known and unknown. A known attacker refers to one whose biometric data has previously been exposed to the system, typically during the training phase. This exposure can influence the system’s learning and adaptation, potentially affecting its ability to accurately identify or reject these attackers later. On the other hand, an unknown attacker is one whose biometric data has never been introduced to the system. This lack of prior exposure means that the system has no learning or adaptation based on this attacker’s data, making it a more realistic and challenging scenario. We expect that unknown attackers represent a higher risk to the system, as they test the system’s ability to authenticate users based purely on learned patterns without prior knowledge of the attacker. While the unknown attacker scenario is realistic and should be considered in evaluations, the known attacker scenario has been used in some papers. Therefore, we provide results for both to demonstrate how unrealistic attacker scenarios can influence outcomes and to encourage researchers to use the unknown attacker scenario.
Figure 2: Schematic depiction outlining the architectural structure of the NeuroIDBench codebase, providing a visual overview of its underlying framework [26]
@@ 2.3 Related Work
Arias et al.(2023) [23] 52,40 $\times$
Fallahi et al.(2023) [27] 41,40
Hernandez et al.(2022) [28] 39 $\times$ $\times$
Bidgoly et al.(2022) [29] 109 $\times$
Maiorana(2021) [30] 45 $\times$ $\times$
Sooriyaarachchi et al.(2020) [31] 20 $\times$ $\times$
Gupta et al.(2020) [32] 20
Nakanishi et al.(2019) [33] 10 $\times$ $\times$
Lin et al.(2018) [11] 179 $\times$ $\times$
Schons et al.(2018) [34] 109 $\times$
Das et al.(2016) [35] 50 $\times$ $\times$
@@
Presently, two primary methods for brainwave authentication are studied in the literature. The first method utilizes shallow classifiers built on time series-based feature extraction techniques, such as Power Spectral Density (PSD) or Autoregressive (AR) models [23, 24, 33, 35]. The second approach involves training specialized deep learning models designed explicitly for brainwave data, aiming to extract features directly associated with user identity [27, 30, 34, 29]. However, traditional feature extraction methods are not sufficiently effective, as they were not originally designed for authentication purposes. Meanwhile, deep learning approaches require a substantial amount of data to develop an adequate feature extractor, which is currently scarce in the field of brainwave authentication. Therefore, it is crucial to determine the most appropriate hyperparameters based on other datasets in the field, which can aid in mitigating these challenges.
@@
Table 1 presents an analysis of 11 papers in the realm of brainwave authentication. Among these, only two papers have made their source code available [27, 32], and merely five have utilized public datasets [27, 32, 23, 29, 34]. Notably, four of these papers employed two common public datasets [36, 37], reflecting the limited data availability, which consequently diminishes the generalizability of the results. On average, these studies included approximately 70 subjects for their evaluation purposes. The subcommunity in brainwave-based authentication clearly require more transparency in the research practice as well as more accessibility in the datasets and source code. To be sure, the brainwave-authentication subcommunity are, in this regard, not alone. For example, Olszewski et al. [38] find that the ”large and long-lived” subcommunity of machine learning security fail to provide data (30%) or source code (40%) even in the papers published at the top-tier security conferences. The authors found out that a full 80% of artifacts are incapable of reproducing the reported results. It is therefore unsurprising that similar problems should plague the new and emerging subcommunity in brainwave-based authentication. Nevertheless, these issues are intensified in brainwave authentication due to the relatively small size of each dataset, heightening the risk of overfitting.
@@
Meanwhile, researchers in the field of Brain-Computer Interface (BCI) encounter similar challenges. Hence, Jayaram and Barachant [39] proposed a benchmark framework [^2] https://github.com/NeuroTechX/moabb for investigating common BCI algorithms across 22 public EEG datasets with over 250 participants. This framework aims to address issues of reproducibility, lack of open-source resources, and result compatibility. However, their study did not cover authentication algorithms. Therefore, it is vital for brainwave authentication research to adopt an open-source approach and to extensively employ existing public datasets. It is crucial to take such steps to enhance the robustness of research in this field and to promote a deeper, more comprehensive exploration of brainwave authentication techniques.
@@
N400 Pijnacker et al. [40] 2017 45 actiCap 32 500 Hz 1
N400 Draschkow et al. [41] 2018 40 BrainAmp, actiChamp 64 1000 Hz 1
N400 Marzecová et al. [42] 2018 18 BrainAmp 59 500 Hz 1
N400 Mantegna et al. [43] 2019 31 BrainAmp, EasyCap 65 500 Hz 1
N400 ERPCORE: N400 [36] 2021 40 Biosemi 30 1024 Hz 1
N400 Hodapp and Rabovky [44] 2021 33 BrainAmp 64 1000 Hz 1
N400 Rabs et al. [45] 2022 38 BrainVision 26 500 Hz 1
N400 Schoknecht et al. [46] 2022 38 ActiCap, ActiChamp 58 500 Hz 1
N400 Toffolo et al. [47] 2022 24 Biosemi 128 512 Hz 1
N400 Lindborg et al. [48] 2022 40 BrainVision 64 2046 Hz 1
N400 COGBCI: Flanker [49] 2023 29 ActiCap, actiChamp 64 512 Hz 3
N400 Stone et al. [50] 2023 64 TMSi Refa 32 512 Hz 1
P300 BrainInvaders12 [51] 2012 25 NeXus-32 16 128 Hz 1
P300 BrainInvaders13a [52] 2013 24 g.GAMMAcap 16 512 Hz 1
P300 BrainInvaders14a [53] 2014 64 g.Sahara 16 512 Hz 1
P300 BrainInvaders14b [54] 2014 37 g.GAMMAcap 32 512 Hz 1
P300 Gao et al. [55] 2014 30 Neuroscan 12 500 Hz 1
P300 BrainInvaders15a [56] 2015 50 g.GAMMAcap 32 512 Hz 1
P300 BrainInvaders15b [57] 2015 44 g.GAMMAcap 32 512 Hz 1
P300 Mouček et al. [58] 2017 250 BrainVision 3 n.a. 1
P300 Hubner et al. [59] 2017 13 BrainAmp DC 31 1000 Hz 1
P300 Sosulski and Tangermann [60] 2019 13 BrainAmp, EasyCap 31 1000 Hz 1
P300 Lee et al. [61] 2019 54 BrainAmp 62 1000 Hz 2
P300 Simões et al. [62] 2020 15 g.tec 8 250 Hz 7
P300 Goncharenko et al. [63] 2020 60 NVX-52 8 500 Hz 1
P300 Chatroudi et al. [64] 2021 24 g.tec 64 1200 Hz 1
P300 Cattan et al. [65] 2021 21 g.USBamp, g.tec 16 512 Hz 1
P300 ERPCORE: P300 [36] 2021 40 Biosemi 30 1024 Hz 1
P300 Won et al. [66] 2022 55 Biosemi 32 512 Hz 1